# Graph Neural Networks for the Offline Nanosatellite Task Scheduling Problem

### v3

@@ 5.4.2 Early-fixing
A heuristic solution can be made by fixing the value assignments based on the most confident biases predicted by SatGNN, similar to what is proposed in the Neural Diving approach by Nair et al. (2020). Compared to the warm-starting approach, instead of merely informing the solver, we modify the instance of the problem and fix the values of the variables that the model is most confident of before calling the solver. More precisely, early-fixing is equivalent (in principle) to adding to the problem constraints of the form
$z_{k}=\hat{z}_{k},\forall k\in\mathcal{C}^{N}_{I},$
which limits $N$ variables to the values predicted by the model. In practice, however, no constraints are added, but the assignment in the existing constraints and the objective replaces the variables. Thus, the resulting sub-problem will contain fewer binary variables and, thus, is expected to be faster to solve.
Note that every feasible solution to the early-fixed problem and the fixed variables is also feasible in the original instance. However, an optimal solution to the sub-problem might not be optimal for the original instance. Even worse, early-fixing the problem might render it infeasible if the model predicts a poor variable assignment. Therefore, this approach of early-fixing is a matheuristic, as it uses mathematical optimization to find heuristic solutions.
@@ 5.4.3 Trust Region
Instead of fixing the variables that the model is most confident of, we may allow a slight deviation from the partial candidate solution. In other words, we can use SatGNN’s prediction to define a trust region (Han et al., 2023). A new constraint is added to the problem of the form
@@
where $\Delta\in\mathbb{N}$ is the parameter that defines the size of the trust region. This limits the solution space to vectors in the vicinity of the partial candidate solution.
@@ 6.2 Feasibility Classification
To further assess the architecture of the SatGNN model, we perform an ablation study on the number of layers in the $NN_{\text{out}}$ component. We repeat the training and evaluation of the SatGNN model in the feasibility classification task across instance sizes for different numbers of layers in the last MLP of our model. More specifically, we experiment with up to 4 layers in $NN_{\text{out}}$, and for each number of layers, we repeat the experiment 5 times with different random initializations. The average performance of those models appears in Figure 6. No significant difference can be observed between the experiments, indicating that the learning primarily occurs before $NN_{\text{out}}$, in the graph convolutional layers. In fact, applying the Wilcoxon signed-rank test (Wilcoxon, 1945) shows no statistically significant ($p$-value$>0.05$) difference between the samples; however, this result should be taken with caution in the face of the size of the samples.
Figure 6: Test set performance of SatGNN with a varying number of layers in $NN_{\text{out}}$. Accuracy is measured as an average over 5 random repetitions of the experiment. Vertical bars indicate the minimum and maximum values observed.
@@ 6.3 Optimal Solution Prediction
We tune the hyperparameters of the models by performing a random search and evaluating the models’ performance on the validation set. More specifically, we optimize the learning rate, number of convolutional layers ($L$), number of hidden features ($d$), graph convolutional operator, and whether or not the convolutions share the parameters. Table 5 shows the ranges and best values found. We refer the reader to our code repository [^5] https://github.com/brunompacheco/sat-gnn for further details, as we have made all our hyperparameter tuning implementation and results available. We underscore that for both models, the SAGE graph convolution is the best choice, as well as sharing the parameters between the two convolutional operators.
@@
Layers ($L$) $\{1,2,3\}$ $2$ $3$
Hidden features ($d$) $\{2^{5},2^{6},2^{7},2^{7}\}$ $2^{6}$ $2^{8}$
[…]
@@ 6.4 SatGNN-based Heuristic
We use the validation set to select the best values for the hyperparameters, aiming to optimize two performance metrics. For once, we select values for the hyperparameters that maximize the expected objective value. The objective values are normalized, i.e., we divide them by the objective value of the best solution known for each instance [^6] We remind the reader that those solutions were computed along with the dataset generation, as detailed in Section 6.1., resulting in a value between 00 (no QoS) and $1$ (max. QoS known for the instance). In parallel, we select the hyperparameters that minimize the time required to find a feasible solution to each problem. The two tunings are performed for the SatGNN model trained with the best solution available and the SatGNN trained with multiple solutions. Table 6 summarizes the best hyperparameters for each model and each heuristic approach, in both evaluations.
Table 6: Best values for partial solution size ($N$) and trust region size ($\Delta$) for each heuristic approach and each SatGNN model, as evaluated on the validation set. Objective corresponds to the values that maximize the objective value of the best solution found. Feasibility indicates the values that minimize the amount of time to find a feasible solution.
@@
SatGNN model Heuristic $N$ $\Delta$ $N$ $\Delta$
@@
Figure 11: Test set performance of the SatGNN-based matheuristics. On the left, we have the distribution of the evaluation metric of interest over the instances of the test set for the multiple approaches, in which the triangle indicates the mean value and the circles indicate outliers. MS indicates that the SatGNN model trained with multiple solutions was used, whereas BS indicates that the model trained solely with the optimal solution was used instead. On the right is the average progress of the lower bound on all test set instances. The objective value is considered relative to the objective of the best solution known, thereby always lying in the unit interval. The heuristics’ hyperparameters $N$ and $\Delta$ are defined upon experiments on the validation set, as presented in Table 6.
To assess the significance of the results presented in Figure 11 through statistical tests. Given that the matheuristic performance data is generated from a shared set of instances, our samples are paired. Furthermore, there is no evidence of normality in the results’ distribution. Therefore, we apply the Wilcoxon signed-rank test (Wilcoxon, 1945), which is a non-parametric version of the t-test for matched pairs. We apply the test pair-wise, comparing each matheuristic approach to every other matheuristic approach, and to the baseline. The results are summarized in the critical difference diagram of Figure 12. Through early-fixing, both SatGNN models provide statistically significant ($p$-value$>0.05$) improvements over the baseline in both performance metrics. However, the results show a significant advantage of the model trained with multiple solutions, providing better solutions (objective value) through early-fixing and trust region, and speeding up the feasibility definition (time to find a feasible solution) through all heuristic approaches.
@@
Figure 12: Critical difference diagram presenting the average test set performance of the SatGNN-based matheuristics (round marker in the axis). A crossbar between two (or more) approaches indicates that their performance (distribution on the test set) was not deemed significantly different ($p$-value$>0.05$) through the paired Wilcoxon signed-rank test (Wilcoxon, 1945).
@@
6.4.1 Partial Solution Size
To evaluate the impact of the value of $N$, the number of variables in the partial solution being fed to the matheuristics, we further evaluate the SatGNN model on the test. More specifically, we vary the value of $N$ for all matheuristics and measure the impact on the test set considering both perspectives (objective value and time to find a feasible solution). For completeness, we also vary $\Delta$, the size of the trust region, to evaluate the full potential of the matheuristic. For conciseness, we consider only the SatGNN model trained with multiple solutions, as it is overall superior, in our experiments, to the model trained solely with the best solution. The results are illustrated in Figure 13.
Figure 13: Performance of the matheuristics for varying values of $N$, the number of binary variables in the partial solution. EF indicates the early-fixing matheuristic, TR the trust region, and WS warm-start. The SatGNN model trained with multiple solutions is used. The evaluation is performed on the test set.
As expected, the number of variables that compose the partial solution is impactful for all approaches, with early-fixing being the one most sensitive to it, and warm-start the less sensitive one. Furthermore, we note a relation between $N$ and $\Delta$ through the peak performance of the trust region method (including early-fixing, which is equivalent to trust region with $\Delta=0$), that is, larger partial solutions seem to require larger trust regions. Intuitively, this result is expected since the partial solutions are based on confidence, and, thus, the larger the partial solution is, the higher the expected probability of it including wrongly predicted variables. Therefore, this results suggests that the SatGNN model is properly trained (its confidence correlates to its performance).
@@ 7 Discussion
The results of the first set of experiments (Sec. 6.2) indicate that the SatGNN can provide accurate feasibility estimations even in challenging scenarios. Three scenarios of increasing difficulty were considered, and the results were quite promising, with over 89% accuracy in all experiments. In the most challenging scenario, SatGNN is evaluated on large ($J\in\{20,22,24\}$) unseen instances, but trained on small instances ($J\in\{9,13,18\}$), and it achieved 94.15% accuracy in feasibility classification. The remarkable accuracy indicates that SatGNN is highly effective at learning the constraints of the ONTS problem. This performance is expected as the feasibility classification task requires the model to learn linear classification boundaries, which are derived from the problem’s linear constraints. In other words, the model learns the indicator function of a convex set (considering the problem’s LP relaxation), for which neural networks are known to be suitable.
@@ Acknowledgments
9 Declaration of Generative AI and AI-assisted technologies in the writing process

### v4

@@ 5.4.2 Early-fixing
Our second matheuristic is similar to the Neural Diving approach proposed by Nair et al. (2020). Instead of merely informing the solver of a partial candidate solution via warm-starting, we modify the instance of the problem and fix the values of the variables that the model is most confident of. More precisely, early-fixing is equivalent to adding constraints of the form
$z_{k}=\hat{z}_{k},\forall k\in\mathcal{C}^{N}_{I}$
to the mathematical program of instance $I$, limiting $N$ variables to the assignments predicted by the model. In practice, however, no constraints are added, but the variables are replaced by the assignment in the model formulation. This is equivalent to a diving heuristic with respect to the branch-and-bound tree of instance $I$. The resulting sub-problem will contain fewer binary variables and, thus, is expected to be faster to solve.
We highlight, however, that while Nair et al. (2020) sample multiple partial solutions from the model output and solve all of them, we sample only one, and with a fixed size. This design choice is based on the fixed time-budget available for the ONTS problem at hand (see Section 3). With a single partial solution of fixed size, we have more control over the amount of binary variables in the program fed to the solver, which correlates to the solving time. In other words, we trade solution quality, which would be expected to be higher if more partial solutions were evaluated, for lower computational cost, since a single problem is solved.
Note that every feasible solution to the early-fixed problem and the fixed variables is also feasible in the original instance. However, an optimal solution to the sub-problem might not be optimal for the original instance. Even worse, early-fixing the problem might render it infeasible if the model’s output result in an infeasible partial assignment. Therefore, this approach of early-fixing is a matheuristic, as it uses mathematical optimization to find heuristic solutions.
@@ 5.4.3 Trust Region
The third approach follows the learning-based heuristic proposed by Han et al. (2023). Instead of fixing the variables that the model is most confident of, we allow a limited deviation from the sampled partial candidate solution. In other words, we define a trust region around the candidate solution.
In practice, a new constraint is added to the mathematical programming formulation of each instance $I$ that limits feasible solutions to a neighborhood around the partial solution predicted by SatGNN. The constraint has the form
@@
where $\Delta\in\mathbb{N}$ is the size of the trust region.
Note that we do not make any distinctions between 1 and 0 assignments, as done in Han et al. (2023). We follow the same confidence-based approach as in the previously proposed methods to select the variables. We do so because there is no a priori bias in optimization problems toward 1 or 0 values. Furthermore, even if the instances in $\mathcal{I}$ exhibit such a bias—as is indeed the case for the ONTS problem—we expect the model to learn it during training, so that it is reflected in the predicted marginal probabilities (and, consequently, in the confidence score).
@@ 6.2 Feasibility Classification
To further assess the architecture of the SatGNN model, we perform an ablation study on the number of layers in the ${\rm NN}_{\text{out}}$ component. We repeat the training and evaluation of the SatGNN model in the feasibility classification task across instance sizes for different numbers of layers in the last MLP of our model. More specifically, we experiment with up to 4 layers in ${\rm NN}_{\text{out}}$, and for each number of layers, we repeat the experiment 5 times with different random initializations. The average performance of those models appears in Figure 6. No significant difference can be observed between the experiments, indicating that the learning primarily occurs before ${\rm NN}_{\text{out}}$, in the graph convolutional layers. In fact, applying the Wilcoxon signed-rank test (Wilcoxon, 1945) shows no statistically significant ($p$-value>0.05>0.05) difference between the samples; however, this result should be taken with caution in the face of the size of the samples.
Figure 6: Test set performance of SatGNN with a varying number of layers in $\rm{NN}_{\text{out}}$. Accuracy is measured as an average over 5 random repetitions of the experiment. Vertical bars indicate the minimum and maximum values observed.
@@ 6.3 Optimal Solution Prediction
We tune the hyperparameters of the models by performing a random search and evaluating the models’ performance on the validation set. More specifically, we optimize the learning rate, number of convolutional layers ($L$), number of hidden features ($d$), graph convolutional operator, and whether or not the convolutions share the parameters. Table 5 shows the ranges and best values found. We provide further details on the hyperparameter tuning, including a sensitivity analysis, in D. Nonetheless, we underscore that the most impactful hyperparameters were the learning rate and the choice of graph convolution. In particular, the choice of the SAGE convolutional operator was strongly correlated with lower validation losses for both the models trained with the best solution and those trained with multiple targets.
@@
Layers ($L$) $\{1,2,3\}$ 22 33
Hidden features ($d$) $\{2^{5},2^{6},2^{7},2^{8}\}$ $2^{6}$ $2^{8}$
Graph convolution $\{\text{GraphConv},\text{SAGE}\}$ SAGE SAGE
Weight tying $\{\text{Yes},\text{No}\}$ Yes Yes
@@ 6.4 SatGNN-based Heuristic
We use the validation set to select the best values for the hyperparameters, aiming to optimize two performance metrics. For once, we select values for the hyperparameters that maximize the expected objective value. The objective values are normalized, i.e., we divide them by the objective value of the best solution known for each instance [^6] We remind the reader that those solutions were computed along with the dataset generation, as detailed in Section 6.1., resulting in a value between 0 (no QoS) and $1$ (max. QoS known for the instance). In parallel, we select the hyperparameters that minimize the time required to find a feasible solution to each problem. The two tunings are performed for the SatGNN model trained with the best solution available and the SatGNN trained with multiple solutions. Table 6 summarizes the best hyperparameters for each model and each heuristic approach, in both evaluations.
Table 6: Best values for partial solution size (NN) and trust region size ($\Delta$) for each heuristic approach and each SatGNN model, as evaluated on the validation set. Objective corresponds to the values that maximize the objective value of the best solution found. Feasibility indicates the values that minimize the amount of time to find a feasible solution.
@@
SatGNN model Heuristic NN $\Delta$ NN $\Delta$
@@
Figure 11: Test set performance of the SatGNN-based matheuristics. On the left, we have the distribution of the evaluation metric of interest over the instances of the test set for the multiple approaches, in which the triangle indicates the mean value and the circles indicate outliers. MS indicates that the SatGNN model trained with multiple solutions was used, whereas BS indicates that the model trained solely with the optimal solution was used instead. On the right is the average progress of the lower bound on all test set instances. The objective value is considered relative to the objective of the best solution known, thereby always lying in the unit interval. The heuristics’ hyperparameters NN and $\Delta$ are defined upon experiments on the validation set, as presented in Table 6.
To assess the significance of the results presented in Figure 11 through statistical tests. Given that the matheuristic performance data is generated from a shared set of instances, our samples are paired. Furthermore, there is no evidence of normality in the results’ distribution. Therefore, we apply the Wilcoxon signed-rank test (Wilcoxon, 1945), which is a non-parametric version of the t-test for matched pairs. We apply the test pair-wise, comparing each matheuristic approach to every other matheuristic approach, and to the baseline. The results are summarized in the critical difference diagram of Figure 12. Through early-fixing, both SatGNN models provide statistically significant ($p$-value>0.05>0.05) improvements over the baseline in both performance metrics. However, the results show a significant advantage of the model trained with multiple solutions, providing better solutions (objective value) through early-fixing and trust region, and speeding up the feasibility definition (time to find a feasible solution) through all heuristic approaches.
@@
Figure 12: Critical difference diagram presenting the average test set performance of the SatGNN-based matheuristics (round marker in the axis). A crossbar between two (or more) approaches indicates that their performance (distribution on the test set) was not deemed significantly different ($p$-value>0.05>0.05) through the paired Wilcoxon signed-rank test (Wilcoxon, 1945).
@@
6.4.1 Sensitivity Analysis
The proposed solution methods share a common parameter: the number $N$ of variables in the partial solution. To evaluate the sensitivity of the results to this parameter, we evaluate the SatGNN model on the test set for varying values of $N$. More specifically, we vary the value of $N$ for all matheuristics and measure the impact on the test set considering both perspectives (objective value and time to find a feasible solution). For conciseness, we consider only the SatGNN model trained with multiple solutions, as it is overall superior, in our experiments, to the model trained solely with the best solution. The results are illustrated in Figure 13.
Beyond $N$, the trust region method has an additional parameter: the size of the trust region, $\Delta$. To evaluate the full potential of this matheuristic, we also consider three different values for $\Delta$ in our sensitivity analysis experiments. These results are also displayed in Figure 13, as the legends indicate.
Figure 13: Sensitivity analysis of the solution methods with respect to the size of the partial solution NN. EF indicates the early-fixing matheuristic, TR the trust region, and WS warm-start. The SatGNN model trained with multiple solutions is used. The evaluation is performed on the test set.
As expected, the number of variables that compose the partial solution is impactful for all approaches, with early-fixing being the one most sensitive to it, and warm-start the less sensitive one. In fact, analyzing the results for the early-fixing approach in detail, we note that $N$ indeed imposes a trade-off between solution quality and speed, as shown in Figure 14, confirming the intuitive interpretation of this parameter discussed in Section 5.4.2. More specifically, increasing the partial solution size seems to speed up the solver, allowing it to find better solutions more quickly, while reducing the final solution quality at the end of the time budget.
Figure 14: Average progress of the lower bound of SatGNN-based early fixing on the test set for varying values of the partial solution size $N$.
Furthermore, we note a relation between $N$ and $\Delta$ through the peak performance of the trust region method (including early-fixing, which is equivalent to trust region with $\Delta=0$). In other words, larger partial solutions seem to require larger trust regions to achieve peak performance. Intuitively, this result is expected since the partial solutions are based on confidence, and, thus, the larger the partial solution is, the higher the expected probability of it including wrongly predicted variables. In turn, it suggests that the SatGNN model is properly trained (its confidence correlates to its performance).
@@ 7 Discussion
The results of the first set of experiments (Sec. 6.2) indicate that SatGNN can provide accurate feasibility estimations even in challenging scenarios. Three scenarios of increasing difficulty were considered, and the results were quite promising, with over 89% accuracy in all experiments. In the most challenging scenario, SatGNN is evaluated on large ($J\in\{20,22,24\}$) unseen instances, but trained on small instances ($J\in\{9,13,18\}$), and it achieved 94.15% accuracy in feasibility classification. The remarkable accuracy indicates that SatGNN is highly effective at learning the constraints of the ONTS problem. This performance is expected as the feasibility classification task requires the model to learn linear classification boundaries, which are derived from the problem’s linear constraints. In other words, the model learns the indicator function of a convex set (considering the problem’s LP relaxation), for which neural networks are known to be suitable.
@@
Finally, we underscore that the computational gains observed in terms of time to find a feasible solution (see, for example, Figure 12(b)) are expected to directly translate to the deployed application. First, because dataset generation and training, which are indeed costly, as discussed above, are performed offline. Then, generating a partial solution with SatGNN takes a negligible amount of time compared to solving the instances. Beyond that, the inference time scales linearly with the size of the graph, which is directly proportional to the instance size (number of variables and constraints). Given that the problem is combinatorial, we can expect the SatGNN inference cost to be totally dominated by the cost of solving the instances for realistic instances.
@@ 8.1 Future Work
Reinforcement learning (RL) has shown promise for learning-based approaches to combinatorial optimization (Jalali Khalil Abadi et al., 2024; O’Malley et al., 2023; Waubert de Puiseau et al., 2022). It has also been explored in settings close to ONTS, such as Earth observation satellite scheduling (He et al., 2022; Jacquet et al., 2024). The possibility of taking sequential decisions makes it easier to incorporate feasibility information into the model. In our case, it could enable the fixing of larger partial solutions without rendering the resulting mathematical program infeasible, reducing the computational cost. It may also reduce dataset-generation effort by learning from interaction rather than labeled optimal solutions. However, developing an RL-based primal heuristic requires careful reward design and an interactive environment, which is substantially different from our supervised learning approach; we therefore leave a full RL solution to future work.
@@ Acknowledgments
Declaration of Generative AI and AI-assisted technologies in the writing process
@@ Appendix B Random instance generation
Appendix C Symbols for SatGNN
The table below provides a summary of the main symbols used in the definition of SatGNN and the related methods.
Symbol Description
   Graph representation
$V_{\text{var}}$ Nodes associated to variables; $|V_{\text{var}}|=n$.
$V_{\text{con}}$ Nodes associated to constraints; $|V_{\text{con}}|=m$.
$E$ Edges between variable nodes and constraint nodes.
$w(e)$ Weight of edge $e\in E$.
$\mathcal{N}(v)$ Neighborhood of node $v\in V$.
   Features and hidden representations
$d_{\text{in}}$ Number of input features for each node.
$\boldsymbol{f}^{v}$ Input feature vector of node $v$.
$F_{\text{var}}$ Matrix of hidden features for variable nodes.
$F_{\text{con}}$ Matrix of input features for constraint nodes.
$L\in\mathbb{N}$ Number of graph convolution layers.
$d\in\mathbb{N}$ Number of hidden features for each node.
$\boldsymbol{h}^{(l)}_{v}$ Hidden feature vector of node $v$ at layer $l\in\{0,\dots,L\}$.
$H^{(l)}_{\text{var}}$ Matrix of hidden features for variable nodes at layer $l$.
$H^{(l)}_{\text{con}}$ Matrix of hidden features for constraint nodes at layer $l$.
$d_{\text{out}}$ Number of output features for each node.
   Fully-connected modules
$\mathrm{NN}_{\text{var}}$ Encoder for variable-node features (produces $\boldsymbol{h}^{(0)}$ on $V_{\text{var}}$).
$\mathrm{NN}_{\text{con}}$ Encoder for constraint-node features (produces $\boldsymbol{h}^{(0)}$ on $V_{\text{con}}$).
$\mathrm{NN}_{\text{out}}$ Output head applied to $\boldsymbol{h}^{(L)}$ (task-dependent output dimension).
   Training
$\mathcal{L}$ Loss function.
$\mathcal{D}$ Dataset of instances and candidate solutions.
$p(\hat{\boldsymbol{z}}\in Z|I)$ True “probability” that the assignment $\hat{\boldsymbol{z}}$ is feasible for instance $I$.
Because there is no stochasticity, it can either be 0 or 1.
$\hat{p}(\hat{\boldsymbol{z}}\in Z|I)$ Estimated “probability” (predicted variable bias) that the assignment
$\hat{\boldsymbol{z}}$ is feasible for instance $I$.
$\hat{p}(\boldsymbol{z}^{*}=1|I)$ Predicted variable biases (towards 1) of an optimal solution for
instance $I$.
   Heuristics
$\kappa(z|I)$ Confidence the model has in each variable bias prediction. Defined as
the maximum over the predicted marginal probabilities.
$\mathcal{C}^{N}_{I}$ set of indices of the $N\in\mathbb{N}$ variables that the model is most confident
with with respect to its prediction for instance $I$.
$\Delta$ Size of the trust region around a predicted partial assignment.
Appendix D Hyperparameter Tuning
We conduct a random search over the hyperparameters listed in Table 5 and evaluate each configuration on the validation set. For each configuration we train SatGNN using the best solution (BS) and the multiple solution (MS) approaches on the training set. The performance is measured as the best validation loss observed in 10 epochs. The best configurations seen for each training method is reported in Table 5.
We evaluate the relevance of tuning each hyperparameter by measuring how sensitive the validation loss of SatGNN is to the different values in search space. The distribution of the validation loss given each hyperparameter choice is illustrated in Figure 15. For each hyperparameter, we test whether validation losses differ across its range using the nonparametric Kruskal–Wallis $H$ test (MacFarland and Yates, 2016). We apply the test independently for the BS and the MS trainings. The only hyperparameters that had a statistically significant ($p\ll 0.05$) impact in the validation loss were the learning rate and the graph convolutional operator, for both training approaches. All other hyperparameters did not achieve significance under this criterion, although the size of the hidden feature vectors $d$ had a strong impact ($H=6.84$ and $p<0.08$) when training with MS.
(a) Best Solution
(b) Multiple Solutions
Figure 15: Validation loss by hyperparameter choice in our hyperparameter tuning experiments.
We refer the reader to our code repository [^7] https://github.com/brunompacheco/sat-gnn for further details, as we have made all our hyperparameter tuning implementation and results available.